OpenStudy (anonymous):

Please, explain me is Z2 a subring of Z6?

OpenStudy (anonymous):

The book says: \(Z_2\) is not even a subset of \(Z_6\). I don't get it.

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (kainui):

Well, what defines a ring?

OpenStudy (anonymous):

a set equipped 2 operation, namely + and * in which for +: associativity , abelian for *: associativity, left/right distribution distribute * over +, that is a(b+c) = ab +ac from both sides (left and right)

OpenStudy (kainui):

Does it have to be closed as well under these operations?

OpenStudy (anonymous):

sure

OpenStudy (kainui):

I am asking because I am partly not sure and partly think that this is the way to the answer you're looking for. It seems to me that in Z2 you will have 1+1=0 and in Z6 1+1=2, but I am going to get out my abstract algebra book because I am curious to figure this out with you now! =D

OpenStudy (anonymous):

My question on it is: if we check the axiom of subring, that is \(Z_2\) has no additive opposite element in it, I agree that \(Z_2\) is not a subring of \(Z_6\) but the reason they give me is \(Z_2\) is not a subset of \(Z_6\) why?

OpenStudy (anonymous):

\(Z_6=\{[0],[1],[2],....,[5]\}\) and \(Z_2=\{[0],[1],[2]\}\) obviously, \(Z_2\subset Z_6\)

OpenStudy (kainui):

I believe that 0, 2, 4 forms a subset of Z6 since every operation on this set is closed.

OpenStudy (kainui):

For example, 0+0=0, 0+2=2, 0+4=4, 2+2=4, 2+4=0, 0*0=0, 0*2=0, 0*4=0, 2*2=4, 2*4=2

OpenStudy (kainui):

You can't do that with Z2 since 0 and 1 won't be closed and form a subset, does that make sense? It's all because 1+1=2 in Z6, otherwise it would work since 0*0=0, 0*1=0, 0+0=0, 0+1=1 all work otherwise.

OpenStudy (xapproachesinfinity):

eh didn't the ring theory, number theory or whatever this is related to Z2 means class with mod 2? right?

OpenStudy (anonymous):

oh, I confused!! what is the definition of a subset of a set?

OpenStudy (anonymous):

@xapproachesinfinity engruent class, not just class

OpenStudy (anonymous):

https://www.google.com/?gws_rd=ssl#q=subset+of+a+set hence, trivial, 

OpenStudy (kainui):

Yeah I don't know the terminology very well, is what I was saying right or are we talking about different problems now?

OpenStudy (kainui):

Oh! Yes look at 24b it is exactly what I am describing, just a different example! =D

OpenStudy (anonymous):

I understand 24b also, :) but 24a hehehe... how???

ganeshie8 (ganeshie8):

Z2 = {0, 1} is a subset of Z6

ganeshie8 (ganeshie8):

Oh you're refering to that link.. idk why it says Z2 is not a subset of Z6

OpenStudy (anonymous):

oh, I put [2] on it :) my bad but it is trivial that Z2 is a subset of Z6, right? why they said so?

OpenStudy (kainui):

I think they meant to write subring? Yeah seems weird, I thought maybe I just didn't know the terminology but now I'm just completely confused since I see what you guys are saying. Good luck!

OpenStudy (anonymous):

Thank you. It is not my assignment. I read and wonder how, just want to clarify it. My problem is on the previous post, not this. :)